xx 1

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11arrayxnx 10Scope: privatex01  INT

11Description: Size of arrays in x direction

11 Range 10  x01  Default: 4

110:*  10

11arrayxny 10Scope: privatex01  INT

11Description: Size of arrays in y direction

11 Range 10  x01  Default: 5

110:*  10

11

11arrayxnz10Scope: privatex01  INT

11Description: Size of arrays in z direction

11 Range 10  x01  Default: 10

110:*  10

11

7 Interfaces

General Implements: testlocalreduce

Grid Variables

7.0.1 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 realxvals 10 x01 compact110
11                 10 realxmaxx3D x01 description11results from real reduction
11                 10 realxmaxx2D x01 dimensions110
11                 10 realxmaxx1D x01 distribution11CONSTANT
11                 10 realxsumx3D x01 group type11SCALAR
11                 10 realxsumx2D x01 timelevels111
11                 10 realxsumx1D x01variable type11REAL
11  intxvals  10  x01  compact110
11                 10  intxmaxx3D  x01 description11results from int reduction
11                 10  intxmaxx2D  x01 dimensions110
11                 10  intxmaxx1D  x01 distribution11CONSTANT
11                 10  intxsumx3D  x01 group type11SCALAR
11                 10  intxsumx2D  x01 timelevels111
11                 10  intxsumx1D  x01variable type11INT

8 Schedule This section lists all the variables which are assigned storage by thorn CactusNumerical/TestLocalReduce. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11realxvals intxvals11x

11x  11x

11

Scheduled Functions

CCTKxPOSTINITIAL x testlocalreducecxmaximum x test out maximum reductions from c x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxPOSTINITIAL x testlocalreducecxsum x test out sum reductions from c x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxPOSTINITIAL x testlocalreducecxavg x test out avg reductions from c x

11x11Language:11c
11 x 11Type: 11function